OpenStudy (demonchild99):

A snow cone is a tasty treat with flavored ice and a spherical bubble gum ball at the bottom: The radius of the cone is 1.25 inches, and its height is 2.75 inches. If the diameter of the bubble gum ball is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice? 4.43 in3 0.07 in3 13.50 in3 4.50 in3

jimthompson5910 (jim_thompson5910):

Do you know the volume formulas for a cone and sphere?

OpenStudy (demonchild99):

...no

jimthompson5910 (jim_thompson5910):

Volume of Cone: \[\Large V = \pi*r^2*h\] Volume of Sphere: \[\Large V = \frac{4}{3}\pi*r^3\]

jimthompson5910 (jim_thompson5910):

where \(\Large \pi \approx 3.14\) (use more decimals of pi to get a more accurate volume)

jimthompson5910 (jim_thompson5910):

So are you able to find the volume of the cone?

OpenStudy (demonchild99):

i would have to multiply 3.14x1.25^2x2.75 right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

that will get you the approximate volume

OpenStudy (demonchild99):

ok 1 sec

OpenStudy (demonchild99):

that would make it c...13.50

jimthompson5910 (jim_thompson5910):

now find the volume of the sphere

jimthompson5910 (jim_thompson5910):

oh my bad

jimthompson5910 (jim_thompson5910):

I wrote the wrong formula for the cone

jimthompson5910 (jim_thompson5910):

The volume of a cone should be \[\Large V = \frac{1}{3}\pi*r^2*h\] so you take 1/3 of 13.50 to get the volume of the cone

OpenStudy (demonchild99):

4.5?

OpenStudy (demonchild99):

4.50*

jimthompson5910 (jim_thompson5910):

roughly, yes, I'd say 4.497 roughly

jimthompson5910 (jim_thompson5910):

how about the sphere?

OpenStudy (demonchild99):

4/3*3.14*.25^3

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what is the value approximately?

OpenStudy (demonchild99):

i got 0.06546875

jimthompson5910 (jim_thompson5910):

looks good

jimthompson5910 (jim_thompson5910):

so you now subtract Answer = (Volume of Cone) - (Volume of sphere)

jimthompson5910 (jim_thompson5910):

the answer you want is the leftover volume after you put the gumball in, which is why we're subtracting

OpenStudy (demonchild99):

approximately 4.43 so A' right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (demonchild99):

thank you so much!

jimthompson5910 (jim_thompson5910):

yw, I'm glad to be of help